OpenStudy (anonymous):

Evaluate cos(31pi/3) I need help solving this, my notes suck.

sam (.sam.):

0.8437

OpenStudy (anonymous):

I think its a reference point or something. my answer sheet has 1/2 , -1/2, \[\sqrt{3}/2\], and -\[\sqrt{3/2}\]

OpenStudy (anonymous):

\[31\pi/3=(30\pi+\pi)/3=5(2\pi)+\pi/3\]

OpenStudy (anonymous):

This means your solution is the same as \[\cos \pi/3\]

OpenStudy (anonymous):

Which should be pretty close to 1/2.

OpenStudy (anonymous):

ok, I thought it was -1/2 , Is there an easier formula to do?

OpenStudy (anonymous):

You need to remove all the multiples of \[2\pi\] that you can, then look at what you have left.

OpenStudy (anonymous):

ok ok, because cos 60 is 1/2 and it's in Q1, so it's positive. Got it!

OpenStudy (anonymous):

Well done.

OpenStudy (anonymous):

Thanks so much.